Zara sells 5% of its fast fashion items at a 20% markdown. If the value of each item is $30 and the cost to make each item is $18, then what is the productivity ratio if Zara sells 1,000 units? (Enter your response rounded to two decimal places.)
Answer #1

Output = Net Revenue

= Revenue from items sold at full price + Revenue from items sold at markdown price

= 1000*(.95*30+.05*30*.8)

= $ 29,700

Input = Total cost

= 1000*18

= $ 18,000

Productivity ratio = Output / Input

= 29700 / 18000

= 1.65